Circular No. 6121

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

PERIODIC COMET WILD 4 (1994v) J. V. Scotti, Lunar and Planetary Laboratory, reports his recovery of

this comet with the Spacewatch telescope at Kitt Peak:

1994 UT  α_{2000}  δ_{2000}  m_{1}
Nov. 9.42947  3^{h}12^{m}17.^{s}07  +20^{o}15′40′′.9  20.7
9.44479  3 12 16.32  +20 15 38.7  20.4
Dec. 9.24248  2 50 28.48  +18 54 43.9
9.24758  2 50 28.30  +18 54 42.7
9.27187  2 50 27.38  +18 54 39.3  20.7
9.27712  2 50 27.19  +18 54 38.1  21.0

The comet’s image is indistinguishable from that of a star, and the indicated correction to the prediction on MPC 22032 is ΔT = −0.3 day.

V635 CASSIOPEIAE H. Mendelson and T. Mazeh, Wise Observatory report: “V635 Cas,

the optical counterpart of the x-ray pulsar 4U 0115+63, is probably in the midst of another optical outburst. Its magnitude was observed in mid-December to be I = 12.2 ± 0.1, similar to the brightness at the peak of previous optical and x-ray outbursts (Mendelson and Mazeh 1991, MNRAS 250, 373). Spectroscopic and x-ray observations are strongly urged.”

SUPERNOVA 1994aj IN ANONYMOUS GALAXY C. Pollas, Observatoire de la Côte d’Azur, reports his discovery of an

apparent supernova (mag ∼ 18) on films obtained on Dec. 12.16 and 13.14 UT. SN 1994aj is located at α = 9^{h}03^{m}46^{s}.12, δ = −10^{o}27′56′′.2 (equinox 1950.0), which is 5′′.2 west and 5′′.1 north of the galaxy’s center. Nothing is visible at this position on the Palomar sky survey prints to mag 20. A nearby star (mag ∼ 17) is at position end figures 44^{s}.50, 28′09′′.1.

N. B. Suntzeff, Cerro Tololo Interamerican Observatory (CTIO), reports that R. Aviles has obtained B and V CCD images with the 0.9-m CTIO telescope on Dec. 14.3 UT of SN 1994aj, confirming that there is a starlike object (with V = 18.32 ± 0.02 and B − V = +0.33 ± 0.03) located 5′′.9 west and 4′′.9 north of the galaxy nucleus.

1994 December 29 Daniel W. E. Green